OpenStudy (anonymous):

4x − 2y if x = −1 and y = −2

sam (.sam.):

4x − 2y if x = −1 and y = −2 =4(-1) − 2(-2) =-4+4 =?

OpenStudy (anonymous):

Thank you so Much < 333333 @.Sam.

sam (.sam.):

welcome :)